OpenStudy (anonymous):

If the probability that it rains tomorrow is 35%, what is the probability that it does NOT rain?

OpenStudy (anonymous):

100% is everything, so just subtract the 35% from the 100%, which is 65%

OpenStudy (anonymous):

65%

OpenStudy (anonymous):

100%- 35% =

OpenStudy (dmezzullo):

ok i should of kept my answer then

OpenStudy (anonymous):

Total Probability = 100 % Total Probability = Probability of Raining + Probability of Not Raining 100 = 35 + x x = 100 - 35 = ??